# Control-Aware Trajectory Predictions for Communication-Efficient Drone Swarm Coordination in Cluttered Environments

### v1

@@ IV-C Performance in testing scenarios
Value of the KKT-informed training procedure. We evaluate the benefits of the KKT-informed training process by comparing VAE+EG+KKT with VAE+EG (Bayesian). We can see that although the prediction performance between the two benchmarks is the same (as shown in Tables II-IV), the control performance of VAE+EG+KKT is significantly better than VAE+EG (Bayesian) (as shown in Tables I, Figure 4, Figure 8, and Figure 10). This is because the KKT-informed training process enables the prediction outcome of VAE+EG+KKT to better align with the embedded QP in DMPC. This can also be reflected by comparing VAE with VAE + KKT which uses the KKT-informed component to enhance the performance of VAE, whereby the only difference is that VAE + KKT integrates a KKT-informed layer during the training procedure. We can see from Tables I, Figure 4, Figure 6, and Figure 9 that using KKT alone can enhance the performance of individual predictors. Notice that due to its high computational burden in forward propagation, we recommend using the KKT component only in the training procedure.
TABLE II: Trajectory prediction accuracy (Channel 1 – $x$-axis)
Time Step
Base-line
EG
VAE
VAE + KKT
VAE + EG (Baye-sian)
VAE + EG + KKT
1
0.41
0.34
0.18
0.09
0.07
0.07
2
0.42
0.30
0.15
0.06
0.05
0.05
3
0.51
0.28
0.14
0.05
0.05
0.05
4
0.56
0.27
0.14
0.05
0.05
0.05
5
0.64
0.26
0.14
0.06
0.05
0.05
6
0.71
0.28
0.15
0.07
0.05
0.05
7
0.77
0.28
0.17
0.08
0.06
0.06
8
0.85
0.30
0.18
0.09
0.07
0.07
9
0.91
0.32
0.19
0.10
0.08
0.08
10
0.97
0.33
0.20
0.11
0.09
0.09
11
1.02
0.35
0.22
0.12
0.10
0.09
12
1.08
0.37
0.24
0.12
0.10
0.10
13
1.14
0.38
0.25
0.13
0.10
0.10
14
1.19
0.39
0.27
0.13
0.11
0.10
15
1.25
0.40
0.28
0.19
0.11
0.10
16
1.32
0.42
0.30
0.16
0.12
0.11
TABLE III: Trajectory prediction accuracy (Channel 2 – $y$-axis)
Time Step
Base-line
EG
VAE
VAE + KKT
VAE + EG (Baye-sian)
VAE + EG + KKT
1
0.15
0.02
0.03
0.01
0.01
0.02
2
0.16
0.01
0.02
0.01
0.01
0.01
3
0.18
0.01
0.02
0.01
0.01
0.01
4
0.21
0.01
0.02
0.01
0.01
0.01
5
0.24
0.01
0.02
0.01
0.01
0.01
6
0.27
0.01
0.02
0.01
0.01
0.01
7
0.31
0.01
0.02
0.01
0.01
0.01
8
0.35
0.01
0.01
0.01
0.01
0.01
9
0.39
0.01
0.01
0.01
0.01
0.01
10
0.44
0.01
0.01
0.01
0.01
0.01
11
0.48
0.01
0.01
0.01
0.01
0.01
12
0.53
0.01
0.01
0.01
0.00
0.00
13
0.57
0.01
0.01
0.01
0.00
0.00
14
0.62
0.01
0.02
0.01
0.00
0.00
15
0.67
0.01
0.02
0.01
0.01
0.01
16
0.71
0.01
0.03
0.01
0.01
0.01
TABLE IV: Trajectory prediction accuracy (Channel 3 – $z$-axis)
Time Step
Base-line
EG
VAE
VAE + KKT
VAE + EG (Baye-sian)
VAE + EG + KKT
1
0.24
0.02
0.03
0.03
0.03
0.03
2
0.24
0.02
0.02
0.02
0.02
0.02
3
0.25
0.03
0.03
0.03
0.02
0.02
4
0.27
0.03
0.03
0.03
0.02
0.02
5
0.29
0.03
0.03
0.03
0.01
0.01
6
0.30
0.02
0.03
0.03
0.01
0.01
7
0.32
0.02
0.02
0.02
0.01
0.01
8
0.34
0.01
0.02
0.01
0.01
0.01
9
0.37
0.01
0.02
0.02
0.01
0.01
10
0.38
0.01
0.02
0.02
0.01
0.01
11
0.41
0.01
0.02
0.01
0.01
0.01
12
0.43
0.01
0.02
0.01
0.01
0.01
13
0.45
0.01
0.02
0.01
0.01
0.01
14
0.47
0.01
0.02
0.01
0.01
0.01
15
0.50
0.01
0.03
0.01
0.02
0.02
16
0.52
0.02
0.04
0.02
0.02
0.02
(a) Distance between UAVs
(b) Minimum distance between obstacles and UAVs
Figure 5: Evolution of distances with Baseline applied.
(a) Distance between UAVs
(b) Minimum distance between obstacles and UAVs
Figure 6: Evolution of distances with VAE applied.
(a) Distance between UAVs
(b) Minimum distance between obstacles and UAVs
Figure 7: Evolution of distances with EG applied.
(a) Distance between UAVs
(b) Minimum distance between obstacles and UAVs
Figure 8: Evolution of distances with VAE+EG (Bayesian) applied.
(a) Distance between UAVs
(b) Minimum distance between obstacles and UAVs
Figure 9: Evolution of distances with VAE + KKT applied.
(a) Distance between UAVs
(b) Minimum distance between obstacles and UAVs
Figure 10: Evolution of distances with VAE+EG+KKT applied.
(a) Distance between UAVs
(b) Minimum distance between obstacles and UAVs
Figure 11: Evolution of distances with DMPC applied.
@@ IV-D Robustness in scenarios with limited communication capabilities
We next demonstrate the robustness of the proposed trajectory prediction algorithm in scenarios with limited communication capabilities by evaluating the sensitivity of the proposed algorithm in these scenarios. Specifically, we evaluate the impact of trajectory compression (due to limited bandwidth, energy requirements, etc.), lower communication frequency, and packet drop on the control performance.
@@
Impact of trajectory compression. We evaluate our trajectory prediction algorithm with various levels of trajectory compression, represented by the dimension of the VAE latent variable $\bm{z}_{k}$ (i.e., the messages exchanged between UAVs). Other parameters remain at the nominal values. Figure 12 shows the safety performance of all scenarios. Table V shows the resulting control performance, where DMPC is added as a reference. As we can see from Table V, the control performance is close to that of DMPC as long as the latent dimension is greater than 24. For lower dimensions, although the control performance drops, the algorithm can still yield safe actions and outperform other benchmarks such as EG, VAE, and VAE+EG (Bayesian) as demonstrated in Table I. This shows that our proposed algorithm can work effectively even with low latent dimensions, suggesting that the algorithm can accommodate relatively low bandwidth.
@@
TABLE V: Sensitivity to the level of trajectory compression (control performance)
Latent Dim.
Safe-Agent Cost
Safe-Obs Cost
Control Effort
Cohesion Cost
Total Cost
6
0
0
165161
109800
274961
18
0
0
145297
97109
242407
24
0
0
159758
51027
210786
30
0
0
151257
53328
204585
DMPC (48)
0
0
141084
67298
208382
Figure 12: Sensitivity to the level of trajectory compression (safety performance)
@@
Impact of lower communication frequencies. We evaluate the robustness of our prediction algorithm in scenarios with various levels of communication frequencies, i.e., 1.25 Hz, 2.5 Hz, and 5 Hz, corresponding to communication every 4,2,1 time steps. Other parameters remain at the nominal values. Figure 13 shows the safety performance of all scenarios. Table VI shows the resulting control performance. As we can see from Table VI, by reducing communication frequency, the control performance drops but is still better than other benchmark algorithms such as EG and VAE. Moreover, we can see that the algorithm can still ensure safety, as the minimum distance between agents and between agents and obstacles is kept higher than the collision threshold. Therefore, we can moderately reduce the communication frequency while ensuring satisfactory control performance.
@@
TABLE VI: Sensitivity to communication frequencies (control performance)
Comm. Freq. (Hz)
Safe-Agent Cost
Safe-Obs Cost
Control Effort
Cohesion Cost
Total Cost
1.25
0
0
140237
156950
297188
2.5
0
0
146800
124174
270974
5
0
0
159758
51027
210786
Figure 13: Sensitivity to communication frequencies (safety performance)
@@
Impact of packet loss. We evaluate the robustness of our prediction algorithm in scenarios with various packet loss probabilities (i.e., 1/2, 1/4, 1/8, 1/16). Other parameters remain at the nominal values. Figure 14 shows the safety performance of all scenarios. Table VII shows the resulting control performance. As we can see from Table VII, by increasing packet loss probabilities, the control performance drops but is still comparable with other benchmark algorithms. Note that with the same amount of communications, packet loss has a more significant impact on the control performance than reducing communication frequencies. This is because reducing communication frequencies can still lead to a well-structured communication graph that evolves according to the dynamics of UAVs, while packet loss may make the communication graph more random. Nevertheless, we can see from Figure 14 that the algorithm can still ensure safety.
@@
TABLE VII: Sensitivity to packet loss probabilities (control performance)
Packet loss pro.
Safe-Agent Cost
Safe-Obs Cost
Control Effort
Cohesion Cost
Total Cost
1/2
0
0
154006
159531
313537
1/4
0
0
149913
106425
256337
1/8
0
0
112363
129328
241692
1/16
0
0
165910
61686
227596
0
0
0
159758
51027
210786
@@
Figure 14: Sensitivity to packet loss probabilities (safety performance)
@@ IV-E Sensitivity to measurement noises
We evaluate the robustness of our prediction algorithm in scenarios with various measurement noises ranging from 0 to 0.032 m. Other parameters remain at the nominal values. Figure 15 shows the safety performance of all scenarios. Table VIII shows the resulting control performance. As we can see from Table VIII, the control performance drops as the magnitude of measurement noises increases. This is expected because as the quality of communication reduces, the measurement becomes more important for UAVs to sense their surroundings. Nevertheless, we can see from Figure 15 that the algorithm can still ensure safety with moderate measurement noises.
TABLE VIII: Sensitivity to measurement noises (control performance)
Noise Level (m)
Safe-Agent Cost
Safe-Obs Cost
Control Effort
Cohesion Cost
Total Cost
0.032
0
0
146342
148548
294889
0.024
0
0
184542
105045
289588
0.016
0
0
140848
117938
258786
0.008
0
0
158898
81059
239957
0
0
0
159758
51027
210786
Figure 15: Sensitivity to measurement noises (safety performance)

### v2

@@ IV-C Performance in testing scenarios
Value of the KKT-informed training procedure. We evaluate the benefits of the KKT-informed training process by comparing VAE+EG+KKT with VAE+EG (Bayesian). We can see that although the prediction performance between the two benchmarks is the same (as shown in Tables II), the control performance of VAE+EG+KKT is significantly better than VAE+EG (Bayesian) (as shown in Tables I,Figure 4, Figure 5k)- 5)p). This is because the KKT-informed training process enables the prediction outcome of VAE+EG+KKT to better align with the embedded QP in DMPC. This can also be reflected by comparing EG with EG+KKT and VAE with VAE+KKT, which use the KKT-informed component to enhance the performance of both EG and VAE. The only difference is that EG+KKT and VAE+KKT integrate the KKT-informed layer during the training procedure. We can see from Tables I, Figure 4, Figure 5e)- 5n) that using KKT alone can enhance the performance of individual predictors. Notice that due to its high computational burden in forward propagation, we recommend using the KKT component only in the training procedure.
Value of communication. We evaluate the benefits of having the communicated message by comparing VAE+EG (Bayesian) and EG, which shows the value of having the reconstructed trajectories from VAE. Our conclusion is similar to the above analysis in that VAE+EG (Bayesian) significantly outperforms EG, as demonstrated in Tables III, as well as in Figure 4 and Figures 5e)-5l).
TABLE II: Trajectory prediction accuracy (meter)
Time Step LIA Y-Net EG VAE EG + KKT VAE + EG (Baye-sian) VAE + KKT VAE + EG + KKT
1 0.49 0.37 0.34 0.18 0.14 0.07 0.06 0.06
2 0.50 0.35 0.30 0.15 0.12 0.05 0.07 0.06
3 0.59 0.37 0.28 0.14 0.12 0.05 0.05 0.05
4 0.65 0.33 0.27 0.14 0.12 0.05 0.05 0.05
5 0.74 0.36 0.26 0.14 0.11 0.05 0.05 0.04
6 0.81 0.37 0.28 0.15 0.13 0.05 0.05 0.05
7 0.88 0.42 0.28 0.17 0.15 0.06 0.07 0.06
8 0.98 0.44 0.30 0.18 0.16 0.07 0.08 0.07
9 1.05 0.45 0.32 0.19 0.17 0.08 0.08 0.08
10 1.13 0.46 0.33 0.20 0.17 0.09 0.08 0.10
11 1.19 0.40 0.35 0.22 0.19 0.10 0.09 0.10
12 1.27 0.44 0.37 0.24 0.22 0.10 0.10 0.11
13 1.35 0.40 0.38 0.25 0.22 0.10 0.11 0.10
14 1.42 0.43 0.39 0.27 0.24 0.11 0.11 0.10
15 1.50 0.44 0.40 0.28 0.25 0.11 0.10 0.10
16 1.58 0.50 0.42 0.30 0.27 0.12 0.12 0.11
((a)) Distance between UAVs with LIA applied.
((b)) Minimum UAV-obstacle distance with LIA applied.
((c)) Distance between UAVs with Y-Net applied.
((d)) Minimum UAV-obstacle distance with Y-Net applied.
((e)) Distance between UAVs with EG applied.
((f)) Minimum UAV-obstacle distance with EG applied.
((g)) Distance between UAVs with VAE applied.
((h)) Minimum UAV-obstacle distance with VAE applied.
((i)) Distance between UAVs with EG + KKT applied.
((j)) Minimum UAV-obstacle distance with EG + KKT applied.
((k)) Distance between UAVs with VAE+EG (Bayesian) applied.
((l)) Minimum UAV-obstacle distance with VAE+EG (Bayesian) applied.
((m)) Distance between UAVs with VAE + KKT applied.
((n)) Minimum UAV-obstacle distance with VAE + KKT applied.
((o)) Distance between UAVs with VAE+EG+KKT applied.
((p)) Minimum UAV-obstacle distance with VAE+EG+KKT applied.
((q)) Distance between UAVs with DMPC applied.
((r)) Minimum UAV-obstacle distance with DMPC applied.
Figure 5: Evolution of distances with different benchmarks applied.
@@ IV-D Robustness in scenarios with limited communication capabilities
We next demonstrate the robustness of the proposed trajectory prediction algorithm in scenarios with limited communication capabilities by evaluating the sensitivity of the proposed algorithm in these scenarios. Specifically, we evaluate the impact of trajectory compression (due to limited bandwidth, energy requirements, etc.), lower communication frequency, and packet loss on the control performance.
@@
Impact of trajectory compression. We evaluate our trajectory prediction algorithm with various levels of trajectory compression, represented by the dimension of the VAE latent variable $\bm{z}_{k}$ (i.e., the messages exchanged between UAVs). Other parameters remain at the nominal values. Table III shows the resulting control performance of all scenarios. Table IV shows the safety performance, where DMPC is added as a reference. As we can see from Table III, the control performance is close to that of DMPC as long as the latent dimension is greater than 24. For lower dimensions, although the control performance drops, the algorithm can still yield safe actions and outperform other benchmarks such as EG, VAE, and VAE+EG (Bayesian) as demonstrated in Table I. This shows that our proposed algorithm can work effectively even with low latent dimensions, suggesting that the algorithm can accommodate relatively low bandwidth.
@@
TABLE III: Sensitivity to the level of trajectory compression (control performance), with costs in units of $10^{5}$
Latent Dim. Safe-Agent Cost Safe-Obs Cost Migra-tion Cost Control Effort Cohesion Cost Total Cost
6 0 0 1.21 1.47 1.30 3.99
18 0 0 1.19 1.06 1.25 3.49
24 0 0 1.18 1.51 0.615 3.30
30 0 0 1.24 1.51 0.533 3.27
DMPC (48) 0 0 1.16 1.41 0.673 3.24
TABLE IV: Sensitivity to the level of trajectory compression (safety performance)
Latent Dim. Min Agent-agent Distance Average Agent-agent Distance Max Agent-agent Distance Min Agent-obstacle Distance
6 0.16 0.70 3.99 0.18
18 0.19 0.65 3.84 0.28
24 0.18 0.66 3.11 0.26
30 0.23 0.70 2.93 0.21
DMPC (48) 0.24 0.73 2.91 0.33
@@
Impact of lower communication frequencies. We evaluate the robustness of our prediction algorithm in scenarios with various levels of communication frequencies, i.e., 1.25 Hz, 2.5 Hz, and 5 Hz, corresponding to communication every 4,2,1 time steps. Other parameters remain at the nominal values. Table V shows the resulting control performance of all scenarios. Table VI shows the safety performance. As we can see from Table V, by reducing communication frequency, the control performance drops but is still better than other benchmark algorithms such as Y-Net and EG. Moreover, we can see that the algorithm can still ensure safety, as the minimum distance between agents and between agents and obstacles is kept higher than the collision threshold. Therefore, we can moderately reduce the communication frequency while ensuring satisfactory control performance.
@@
TABLE V: Sensitivity to communication frequencies (control performance), with costs in units of $10^{5}$
Comm. Freq. (Hz) Safe-Agent Cost Safe-Obs Cost Migra-tion Cost Control Effort Cohesion Cost Total Cost
1.25 0 0 1.77 0.936 1.82 4.52
2.5 0 0 1.84 1.92 0.112 3.87
5 0 0 1.18 1.51 0.615 3.30
TABLE VI: Sensitivity to communication frequencies (safety performance)
Comm. Freq. (Hz) Min Agent-agent Distance Average Agent-agent Distance Max Agent-agent Distance Min Agent-obstacle Distance
1.25 0.25 0.73 2.86 0.30
2.5 0.29 0.63 2.01 0.31
5 0.18 0.66 3.11 0.26
@@
Impact of packet loss. We evaluate the robustness of our prediction algorithm in scenarios with various packet loss probabilities (i.e., 1/2, 1/4, 1/8, 1/16). Other parameters remain at the nominal values. Table VII shows the resulting control performance of all scenarios. Table VIII shows the safety performance. As we can see from Table VII, by increasing packet loss probabilities, the control performance drops but is still comparable with other benchmark algorithms. Note that with the same amount of communications, packet loss has a more significant impact on the control performance than reducing communication frequencies. This is because reducing communication frequencies can still lead to a well-structured communication graph that evolves according to the dynamics of UAVs, while packet loss may make the communication graph more random. Nevertheless, we can see from Table VIII that the algorithm can still ensure safety.
@@
TABLE VII: Sensitivity to packet loss probabilities (control performance), with costs in units of $10^{5}$
Packet Loss Pro. Safe-Agent Cost Safe-Obs Cost Migra-tion Cost Control Effort Cohesion Cost Total Cost
1/2 0 0 1.24 1.39 1.40 4.02
1/4 0 0 1.28 1.45 1.10 3.84
1/8 0 0 1.22 1.39 1.12 3.73
1/16 0 0 1.26 1.37 0.784 3.42
0 0 0 1.18 1.51 0.615 3.30
@@
TABLE VIII: Sensitivity to packet loss probabilities (safety performance)
Packet loss pro. Min Agent-agent Distance Average Agent-agent Distance Max Agent-agent Distance Min Agent-obstacle Distance
1/2 0.27 0.68 3.83 0.33
1/4 0.23 0.73 3.73 0.10
1/8 0.23 0.67 3.33 0.29
1/16 0.19 0.72 3.46 0.19
0 0.18 0.66 3.11 0.26
@@ IV-E Sensitivity to measurement noises
We evaluate the robustness of our prediction algorithm in scenarios with various measurement noises ranging from 0 to 0.032 m. Other parameters remain at the nominal values. Table IX shows the resulting control performance of all scenarios. Table X shows the safety performance. As we can see from Table IX, the control performance drops as the magnitude of measurement noises increases. This is expected because as the quality of communication reduces, the measurement becomes more important for UAVs to sense their surroundings. Nevertheless, we can see from Table X that the algorithm can still ensure safety with moderate measurement noises.
TABLE IX: Sensitivity to measurement noises (control performance), with costs in units of $10^{5}$
Noise Level (m) Safe-Agent Cost Safe-Obs Cost Migra-tion Cost Control Effort Cohesion Cost Total Cost
0.032 0 0 1.10 1.31 1.67 4.08
0.024 0 0 1.12 1.35 1.52 3.98
0.016 0 0 1.10 1.29 1.33 3.72
0.008 0 0 1.11 1.35 1.16 3.61
0.004 0 0 1.18 1.51 0.615 3.30
TABLE X: Sensitivity to measurement noises (safety performance)
Noise Level (m) Min Agent-agent Distance Average Agent-agent Distance Max Agent-agent Distance Min Agent-obstacle Distance
0.032 0.14 0.67 3.56 0.31
0.024 0.17 0.74 3.37 0.12
0.016 0.23 0.71 3.62 0.31
0.008 0.19 0.74 3.40 0.22
0.004 0.18 0.66 3.11 0.26